OpenStudy (anonymous):

what is the simplelest form of the following expression? 6c^2+2.5d-d+2c^2-3d a) 4c^2-1.5d b)6c^2+0.5d c)8c^2-1.5d d)8c^2+1.5d

OpenStudy (anonymous):

6c^2 + 2.5d - d + 2c^2 - 3d combine like terms 6c^2 + 2c^2 = 8c^2 2.5d - 1d - 3d = 2.5d - 4d = - 1.5d now put them together 8c^2 - 1.5d any questions ?